# Reinforcement Learning with Foundation Priors: Let the Embodied Agent Efficiently Learn on Its Own

### v3

@@ 4 Experiments
In this section, we present detailed evaluations of the Foundation-guided Actor-Critic (FAC) on robotics manipulation tasks, encompassing both real-world and simulated environments. Additionally, we explore the impact of foundation prior knowledge through ablations in simulation, specifically focusing on sample efficiency and robustness. Our experiments are designed to address the following questions: (a) How sample efficient is the FAC algorithm, in Sec. 4.1 and Sec. 4.2; (b) What is the individual significance of the three foundation priors, in Sec. 4.3; (c) How does the quality of the foundation model influence the performance of FAC, in Section 4.3. More Ablations are included in App. A.5 , and the running time analysis of foundation priors is in App. A.4. Demo real-world videos and generated videos from the diffusion model are attached in the supplementary.
@@ 4.1 Manipulation tasks on Real Robots
Experimental Setup. We established a real-world tabletop environment using a Franka Emika Panda robot, which is equipped with a 7-degree-of-freedom (DoF) arm and a 1-DoF parallel jaw gripper. Observations are obtained from RGB images captured by both a fixed external camera and a wrist-mounted camera on the robot. We design five dexterous manipulation tasks to evaluate the performance of FAC, as illustrated in Fig. 3. The agent undergoes an hour of learning with the environments for these tasks, except for the "Pick Place" task, which involves a training period of 30 minutes. The numbers of collected real-world trajectories of the tasks Pick Place, Open Door, Watering Plants, Unscrew Bottle Cap, and Play Golf are 40, 75, 60, 50, and 115 respectively. These tasks are designed to exhibit the necessity for leveraging prior knowledge to accurately manipulate the objects, which can be hard for vanilla RL. For each task, the starting position of the Franka Arm remains fixed, while the positions of objects vary within a predefined range across different trajectories. And we evaluate them across 10 distinct trajectories under small variations in the environment. Further training details are provided in App. A.2.
Acquring Foundation Prior. (1) Prior success-reward. We employ GPT-4V to automatically determine the success-reward. For efficiency, each task maintains a consistent episodic length, with only the final observation being evaluated for success determination. To assess the accuracy of GPT-4V, we sampled 20 trajectories for each task. Remarkably, the results demonstrate GPT-4V’s robustness in correctly identifying success states, with no recorded false negatives or false positives. The exception is the Watering Plants task, where the false-positive rate was 25%. This is attributed to the task’s specific requirement: successful watering occurs only when the spout orients towards the plants horizontally, which is more challenging. Overall, GPT-4V can serve as an effective proxy model for acquiring success-reward prior knowledge. (2) Prior policy. We explore GPT-4V to make code generations for the initial state. Following previous works [40, 41, 34], we provide some primitive low-level skills, such as move_to, grasp, release, rotate_anticlockwise, etc, including the usage and the parameters of the functions. It is an executable code that takes as input the current robot state and outputs an action based on the configuration and the current state. The code policy can output a prior action based on the robot’s state. Therefore, we can take the KL loss between the policy and prior policy actions on non-initial states (3) Prior value. We load the VIP [24] model, which is a universal value model trained on some internet-scale datasets. We input the goal observation and the current observation to infer the value of the state. Demonstrations of the prompts, the generated policy, and the setting details of each prior are available in App. A.3.1.
@@
Figure 5: Prior policy attempts to open the door without a successful grasp, whereas FAC persistently tries to secure the handle before pulling back the arm through training.
Baselines and Performance Analysis in Real. On real robots, we compare our method to the following baselines: (1) Vanilla RL, utilizing DrQ-v2 [11] with manually designed rewards; (2) Policy Prior, which executes the code policy generated from GPT-4V. The results are in Tab. 1. We find that learning from scratch under the vanilla RL fails in all tasks in one hour of training, because it experiences large amounts of aimless exploration. The code policy prior can achieve 22% success rates on average because the VLM has the commonsense knowledge of how to solve tasks, which can provide some reasonable information for the agents. Significantly, under the guidance of the foundation priors, FAC has an impressive average success rate of 86% across all the tasks.
Furthermore, the learned policy has the ability to adapt and refine its approach based on the prior policy, so as to successfully accomplish tasks, as exemplified in Fig. 5. Initially, the prior policy struggles with grasping the door handle, attempting to open the door without a successful grasp. In contrast, FAC persistently tries to secure the handle before pulling back the arm, which significantly enhances performance. The remarkable performance of FAC proves the efficiency of our proposed framework. It indicates the substantial potential to leverage abundant prior knowledge for embodied agents under the RLFP framework.
@@ 4.2 Manipulation tasks in Simulations
Acquiring Foundation Prior. (1) Prior success-reward. There are few foundation models to distinguish the success behavior in embodied AI, and we find that GPT-4V usually fails in simulation images. So we choose the ground-truth 0-1 success reward in simulation. We also distill a noisy success-reward model to ablate, in Sec. 4.3. (2) Prior policy. To demonstrate that the prior model can be agnostic to the forms, we choose diffusion foundation policy prior. We follow the pipeline of UniPi [25], which generates videos from diffusion models and infers actions from an inverse dynamics model. To make it more efficient, we offline distill a policy model by generating videos from diffusion models and inferred actions from the collected videos. We choose an open-sourced video diffusion model Seer [26], which predicts videos conditioned on images and language instructions. Ideally, the model can be plugged in without in-domain fine-tuning. However, we find the current open-source video prediction models fail to generate reasonable videos in the simulator. Consequently, we fine-tune the Seer with 10 example videos from each task. Compared to UniPi which fine-tunes with 200k videos, the videos generated by our model are much more noisy, attached in supplementary. (3) Prior value. We choose the model VIP [24], and keep the same usage as the experiments on real robots. Implementation details of the diffusion policy are in App. A.3.2.
@@
Performance Analysis in Meta-World. We compare the performance of our method with the above baselines on 8 tasks in Meta-World with 1M frames. Our proposed FAC achieves 100% success rates in all the tasks. 7/8 of them require fewer than 100k frames (about 1-hour training time). For the hard task bin-picking, FAC requires fewer than 400k frames. In contrast, the baseline methods can not achieve 100% success rates on most of tasks. As illustrated in Fig. 6, the sample efficiency and the success rates of FAC are much superior compared to the baseline methods. DrQ-v2 is able to complete some tasks but learns much slower compared to FAC. Moreover, we collect 10 success demos by the diffusion-based prior policy, and warm up the actor by the success demos before training. For the prior policy with 0 success rate, we collect 10 demos. The results are shown in Fig. 7. The improvement of DrQ-v2 with warmup by success demos is evident, but it still lags behind FAC generally. This is because FAC benefits from policy prior guidance during the training process.
R3M and VIP backbones inject the visual representation prior knowledge into the RL process, but their performance is even worse than DrQ-v2. We hypothesize that it might be caused by the pre-trained model having lost plasticity [66]. Since the UniPi and the distilled foundation prior policy do not involve interaction with the environments in training, they are represented as two horizontal lines in Fig. 6. UniPi outperforms the distilled prior in most environments, since the distilled prior is learned from UniPi. However, UniPi is still far inferior compared to FAC.
@@ 4.3 Ablation Study in Simulations
Ablation of Each Foundation Prior. To assess the significance of the three foundation prior, we conduct experiments by individually omitting each prior and comparing the results with the full method. Additionally, we included an ablation study focused on success trajectory imitation, a form of policy regularization. For convenience, we note the success buffer as the usage of success trajectory imitation in the result figure. Figure 8 (a) shows the four ablations.
We find that the reward prior is the most important, without which the performance over all the tasks drops a lot or even fails. The reason is that, without the 0-1 success signals, the reward function is reduced to the shaping reward, where any policy is equivalent under this reward. Without the policy prior, the agent fails on some hard tasks, such as bin-picking and door-open. It also converges much more slowly on most of tasks, such as door-unlock. Without the value prior or the success buffer, the sample efficiency would drop, especially for the hard task bin-picking. This is because, under noisy policy prior, the shaping rewards inferred from value prior can guide the policy towards states of higher values more effectively. Success trajectory imitation learns from successful trials, thereby enhancing sample efficiency. Generally, employing all foundation priors is the optimal strategy.
@@
FAC with Various Quality of Foundation Priors. As shown in previous ablations, without value prior knowledge, the sample efficiency of FAC will drop by a degree. It indicates that a better value foundation prior can further boost performance. Here, we investigate how FAC is robust to the quality of the policy and success reward foundation priors. Firstly, we design several noisier policy priors. We discretize each action dimension generated from the distilled policy model into only three values $\{-1,0,+1\}$. This makes the policy prior only contain rough directional information. We name this prior as discritized policy. To generate even noisier prior, we replace the discretized actions with uniform noise at 20% and 50% probability. As shown in Figure 8 (b1), we find that the discretized policy prior (in the blue curve) performs similarly to the original policy prior (in the green curve), except for bin-picking and door-unlock, which requires more frames to achieve 100% success rates. Adding extra noises decreases the performance further. However, even using the discretized policy with 50% noise, FAC can still reach 100% success rates in many environments. To further investigate the robustness, we apply systematically wrong policy prior, which has a 20% or 50% probability to make the prior action inverse (-1 changes to 1) in Figure 8 (b2). We find that FAC works well under the 20% ratio of the wrong direction, but the success rate drops a lot under the 50% ratio, which is full of misleading information. The results under the 20% ratio indicate that our method has good robustness to systematical noises.
Regarding the quality of the success-reward prior, we distill a proxy model with 50k offline data from the replay buffer, in total for all 8 tasks. The proxy model takes images as input and is conditioned on the task embeddings (multi-task), which has a 1.7% false positive error and 9.9% false negative error on the evaluation datasets. We replace the oracle 0-1 success reward with the predicted success reward. As depicted in Fig. 8 (c), compared to the oracle success-reward, FAC under the 50k-images-distilled model has a limited performance drop in the tasks generally. Moreover, it can achieve much superior performance than the FAC w.o. success reward. Consequently, the proposed FAC can work well under the noisy success-reward foundation prior. In summary, our ablation studies consistently demonstrate the resilience of FAC to variations in the quality of foundation priors. The higher the quality of the prior, the more sample-efficient FAC becomes.
@@ 5 Discussion
In this paper, we introduce a novel framework, termed Reinforcement Learning from Foundation Priors (RLFP), which leverages policy, value, and success-reward prior knowledge for RL. Additionally, we also detail the implementation of this concept within actor-critic methods, introducing the Foundation-guided Actor-Critic (FAC) approach. Extensive experiments on real robots and simulated environments demonstrate the effectiveness of our framework in enabling efficient autonomous learning by agents.
Limitation This work does have its limitations. It still relies on human engineering in designing the low-level skills and the prompts, such as providing the code policy demos and task descriptions. For the poor diffusion foundation models, we need to fine-tune with a few in-domain data to provide reasonable prior knowledge. However, as the performance of foundation models continues to improve, in the future, we may only need to provide very simple task instructions as prompts, without the need for fine-tuning the models.
Looking ahead, there are two dimensions for future exploration of this work. Firstly, it is imperative to construct accurate and broadly applicable foundation priors. Additionally, it is promising to introduce more abundant prior knowledge for the RLFP framework. For instance, humans can predict the future states. Such prediction prior knowledge can be extracted from the dynamic foundation models, which can be potentially more effective for policy learning.
@@ A.4.1 Analysis of Extra Running Clock Time in FAC
Tab. 4 and 5 are the results of real robots and simulations concerning clock time during training. Here, Action Move means the process of taking actions by the robotic arm. The Action Inference means the inference time of the learned policy model. Notably, in the simulation, we offline distill a prior policy from the diffusion model due to the heavy time-cost of video generation (in Sec. LABEL:subsec:method:acquire_prior). Thus, the clock time of the Diffusion Policy Prior for video generation is not included in the training time, although it is much more time-consuming than the other operations in simulation.

### v4

@@ 4 Experiments
In this section, we provide detailed evaluations of the Foundation-guided Actor-Critic (FAC) on robotics manipulation tasks in both real-world and simulated environments. We also examine the impact of foundation prior knowledge through ablations in simulation, focusing on sample efficiency and robustness. Our experiments aim to answer the following questions: (a) How sample-efficient is FAC (Sec. 4.1 and Sec. 4.2); (b) What is the significance of each foundation prior (Sec. 4.3); (c) How does the quality of the foundation model affect FAC’s performance (Sec. 4.3). Additional ablations and running time analysis are in App. A.5. Demo videos are included in the supplementary.
@@ 4.1 Manipulation tasks on Real Robots
Experimental Setup. We set up a real-world tabletop environment using a Franka Emika Panda robot with a 7-DoF arm and a 1-DoF parallel jaw gripper. Observations are collected from RGB images captured by a fixed external camera and a wrist-mounted camera. We designed five dexterous manipulation tasks to evaluate FAC, shown in Fig. 3. The agent learns for one hour per task, except for "Pick Place," which trains for 30 minutes. The number of real-world trajectories collected for Pick Place, Open Door, Watering Plants, Unscrew Bottle Cap, and Play Golf are 40, 75, 60, 50, and 115, respectively. These tasks highlight the need for leveraging prior knowledge for accurate manipulation, which can be challenging for standard RL. The Franka Arm’s starting position is fixed for each task, but object positions vary within a predefined range. We evaluate each task across 10 trajectories with slight environment variations. Additional training details are in App. A.2.
Acquring Foundation Prior. (1) Prior Success-Reward: We use GPT-4V to determine success-rewards, evaluating only the final observation of each task. In a test of 20 trajectories per task, GPT-4V demonstrated high accuracy, with no false negatives or positives, except in the Watering Plants task, which had a 25% false-positive rate due to the challenge of correctly orienting the spout. Overall, GPT-4V is an effective proxy for success-reward prior knowledge. (2) Prior Policy: We use GPT-4V to generate code for initial states, building on previous work [40, 41, 34] by providing primitive low-level skills like move_to, grasp, release, and rotate_anticlockwise. The generated code, based on the robot’s state, outputs prior actions, allowing us to apply KL loss between the policy and prior actions in non-initial states. (3) Prior Value: We use the VIP model [24], a universal value model trained on large-scale datasets. By inputting goal and current observations, we infer the value of states. Detailed prompts, generated policies, and prior settings are in App. A.3.1.
@@
Figure 5: Prior policy attempts to open the door without a successful grasp, whereas FAC persistently tries to secure the handle before pulling back the arm.
Baselines and Performance Analysis in Real. We compared our method to two baselines on real robots: (1) Vanilla RL using DrQ-v2 [11] with manually designed rewards, and (2) Policy Prior using the GPT-4V-generated code policy. Results are shown in Tab. 1. Vanilla RL failed in all tasks after one hour of training due to excessive exploration. The code policy prior achieved an average success rate of 22%, benefiting from VLM’s commonsense knowledge. Notably, FAC, guided by foundation priors, achieved an impressive average success rate of 86% across all tasks.
The learned policy in FAC adapts and refines its approach based on the prior policy to successfully complete tasks, as shown in Fig. 5. While the prior policy struggles with grasping the door handle, FAC persistently secures the handle before pulling, significantly improving performance. FAC’s impressive results demonstrate the efficiency of our framework and highlight the potential of leveraging abundant prior knowledge for embodied agents using the RLFP framework.
@@ 4.2 Manipulation tasks in Simulations
Acquiring Foundation Prior. (1) Prior Success-Reward: Few foundation models can distinguish success behaviors in embodied AI, and GPT-4V often fails with simulation images. Therefore, we use the ground-truth 0-1 success reward in simulation and distill a noisy success-reward model for ablation in Sec. 4.3. (2) Prior Policy: To show that the prior model can be agnostic to form, we use a diffusion-based policy prior, following the UniPi [25] pipeline, which generates videos using diffusion models and infers actions from an inverse dynamics model. For efficiency, we offline distill a policy model from videos generated by the open-source Seer [26] model, which predicts videos conditioned on images and language instructions. While in-domain fine-tuning is ideally unnecessary, current models fail in the simulator. Thus, we fine-tuned Seer with 10 example videos per task, though these videos are much noisier than the 200k used by UniPi, as shown in the supplementary materials. (3) Prior Value: We use the VIP model [24] with the same setup as in the real robot experiments. Implementation details of the diffusion policy are in App. A.3.2.
@@
Performance Analysis in Meta-World. We compared our method to baselines on 8 Meta-World tasks with 1M frames. FAC achieved 100% success rates in all tasks, with 7/8 requiring fewer than 100k frames (about 1 hour of training), and the harder bin-picking task needing under 400k frames. In contrast, baseline methods fail to reach 100% success on most tasks. As shown in Fig. 6, FAC significantly outperforms baselines in both sample efficiency and success rates. While DrQ-v2 can complete some tasks, it learns much slower than FAC. Additionally, we warm up the actor with 10 success demos from the diffusion-based prior policy before training. Although the improvement in DrQ-v2 with warmup is noticeable (Fig. 7), it falls behind FAC, which benefits from policy prior.
R3M and VIP backbones provide visual representation prior knowledge for RL, but perform worse than DrQ-v2, likely due to the loss of plasticity in pre-trained models [66]. As UniPi and the distilled foundation prior policy lack environment interaction during training, they are represented as horizontal lines in Fig. 6. While UniPi outperforms the distilled prior in most environments, since the latter is learned from UniPi, both are still far inferior to FAC.
@@ 4.3 Ablation Study in Simulations
Ablation of Each Foundation Prior. To assess the importance of each foundation prior, we conducted experiments by removing them individually and comparing the results to the full method, along with an ablation study on the success buffer (Fig. 8 (a)). We found the reward prior to be the most crucial. Without it, performance drops significantly or fails, as the reward function reduces to shaping rewards, making policies indistinguishable. Without the policy prior, the agent struggles with difficult tasks like bin-picking and door-opening, and convergence slows on most tasks. Removing the value prior or success buffer reduces sample efficiency, especially for bin-picking, as the value prior helps guide policies under noisy policy priors. Success trajectory imitation further improves sample efficiency. Overall, using all foundation priors yields the best results.
@@
FAC with Various Quality of Foundation Priors. As previous ablations show, without value prior knowledge, FAC’s sample efficiency decreases, suggesting that a better value prior can further boost performance. We now explore FAC’s robustness to the quality of policy and success reward priors. First, we create noisier policy priors by discretizing each action dimension into ${-1,0,+1}$, providing only rough directional information, which we call the discretized policy. We also add uniform noise to the discretized actions at 20% and 50% probabilities. As shown in Fig. 8 (b1), the discretized policy (blue curve) performs similarly to the original (green curve), except for harder tasks like bin-picking and door-unlocking, which take more frames to reach 100% success. Adding noise further reduces performance, but even with 50% noise, FAC still achieves 100% success in many environments. We also tested robustness with systematically wrong policy priors, where actions have a 20% or 50% chance of being inverted (e.g., -1 to 1) (Fig. 8 (b2)). FAC performs well with 20% wrong direction but struggles at 50%, where misleading information is abundant. This shows FAC’s robustness to moderate systematic noise.
To assess the impact of success-reward prior quality, we distilled a proxy model using 50k offline data from the replay buffer across all 8 tasks. The model, conditioned on task embeddings, has a 1.7% false positive and 9.9% false negative error on the evaluation datasets. We replaced the oracle 0-1 success reward with the predicted reward. As shown in Fig. 8 (c), FAC with the 50k-image proxy model experiences only a limited performance drop compared to the oracle reward and outperforms FAC without success rewards. This demonstrates that FAC performs well even with a noisy success-reward prior. In conclusion, our ablation studies show that FAC is resilient to variations in the quality of foundation priors. The higher the quality, the more sample-efficient FAC becomes.
@@ 5 Discussion
In this paper, we introduce a novel framework, termed Reinforcement Learning from Foundation Priors (RLFP), which leverages policy, value, and success-reward prior knowledge for RL. Additionally, we also detail the implementation of this concept within actor-critic methods, introducing the Foundation-guided Actor-Critic (FAC) approach. Extensive experiments on real robots and simulated environments demonstrate the effectiveness of RLFP in efficient autonomous learning.
Limitation and Future Work This work has limitations, such as relying on human engineering for designing low-level skills, prompts, and code policy demos. Additionally, we fine-tune diffusion models with in-domain data for better prior knowledge. However, as foundation models improve, future methods may only require simple task instructions without fine-tuning. Future exploration can focus on building more accurate and broadly applicable foundation priors and incorporating richer priors into the RLFP framework. For example, humans can predict future states, and integrating the predictive knowledge from dynamic foundation models could enhance policy learning.
Acknowledgments
This work is supported by the Ministry of Science and Technology of the People´s Republic of China, the 2030 Innovation Megaprojects "Program on New Generation Artificial Intelligence" (Grant No. 2021AAA0150000). This work is also supported by the National Key R&D Program of China (2022ZD0161700).
@@ A.4.1 Analysis of Extra Running Clock Time in FAC
Tab. 4 and 5 are the results of real robots and simulations concerning clock time during training. Here, Action Move means the process of taking actions by the robotic arm. The Action Inference means the inference time of the learned policy model. Notably, in the simulation, we offline distill a prior policy from the diffusion model due to the heavy time-cost of video generation (in Sec. 3.3). Thus, the clock time of the Diffusion Policy Prior for video generation is not included in the training time, although it is much more time-consuming than the other operations in simulation.